OpenStudy (anonymous):

A baseball player with mass m = 79 kg, sliding into second base, is retarded by a frictional force of magnitude 470 N. What is the coefficient of kinetic friction µ between the player and the ground? [Halliday, Resnick and Walker 2003, 112].

OpenStudy (lgbasallote):

\[f = \mu k \cdot F_N\] to solve for \(\mu k\) dvide both sides by F_N \[\mu k = \frac{f}{F_N}\] F_N is just mg (agree?) \[\mu k = \frac{f}{mg}\] so now substitute \[\mu k = \frac{470}{(79)(9.8)}\] use a calculator to solve does that help?

OpenStudy (anonymous):

ok i will...thank you so much :)

OpenStudy (lgbasallote):

welcome